OpenStudy (anonymous):

Solve for x: 5x – 8 = x + 4

mathslover (mathslover):

\[\huge{5x-x=4+8}\] \[\huge{4x=12}\]\[\huge{x=\frac{12}{4}=3}\]

OpenStudy (anonymous):

^^ Your smart xD

OpenStudy (anonymous):

DO you understand what @mathsolver did?

OpenStudy (anonymous):

Sortaa .

OpenStudy (anonymous):

Alright, well you had your equation with that lovely picture up there that I'm not going to reproduce. Lets look at the first half of the equation 5x-x.

OpenStudy (anonymous):

What its basically saying is 5x-1x. Then you just ignore the x and subtract the numbers. So 5-1=4. Then you add the x back. 5x-1x=4x

OpenStudy (anonymous):

Where did you get the 1 ?

OpenStudy (anonymous):

I just put it there so you could see the subtraction in action. Cause all a number infront of x does is multiply it, so (4)(x)=4x. (1)(x)=x

OpenStudy (anonymous):

1x=x

OpenStudy (anonymous):

just like (1)(132)=132

OpenStudy (anonymous):

Got it?